Facts For the year ended December 31, YR01 Target Inc. had net income of $10,000. At...

Facts

For the year ended December 31, YR01 Target Inc. had net income of $10,000. At this date

Target had 1,000 common shares issued and outstanding (par value =$1 per share). The number of common shares issued and outstanding did not change during YR01. Also, Target Inc. has no outstanding convertible securities. Target Inc. issued its December 31, YR01 annual report to shareholders on February 1, YR02. On January 15, YR02 the Board of Directors declared a 10% stock dividend to be distributed on January 20, YR02. At the date of declaration, the market price of the common stock was $10 per share.

Question

Does the stock dividend declared and distributed in early YR02 affect the EPS disclosure that

will appear in the December 31, YR01 annual report?

1. Provide a brief written description of the proper treatment for the January 15, YR02 stock

dividend with respect to the December 31, YR01 EPS disclosure.

2. Compute the EPS disclosure that will appear in the December 31, YR01 annual report.

3. Identify the specific paragraph of the FASB Codification which addresses this issue and

submit a printout of this paragraph with your solution.
Answer #1
  1. On December 31, YR01 Target INC. had 1,000 common shares issued and outstanding. However, on January 15, YR02, there was 10% stock dividend to be declared. Because the stock dividend occurred after the end of accounting period but before the financial statement was published February 1, YR02, the weighted average shares outstanding must be adjusted for the effects of the stock dividend.

2. Net Income= $10,000

Weighted Average Shares Outstanding= (1,000 x 10%) + 1,000 = 1,100 shares.

EPS= 10,000/1,100= $9.09

  1. The specific paragraph of the FASB Codification which addresses this issue:

FASB Codification “Restatement EPS Data for Stock Dividend and Stock Split” 20-10-55-12 stated that “If the number of common shares outstanding increases as a result of a stock dividend or stock split (see Subtopic 505-20) or decreases as a result of a reverse stock split, the computations of basic and diluted EPS shall be adjusted retroactively for all periods presented to reflect that change in capital structure. If changes in common stock resulting from stock dividends, stock splits, or reverse stock splits occur after the close of the period but before the financial statements are issued or are available to be issued (as discussed in Section 855-10-25), the per-share computations for those and any prior-period financial statements presented shall be based on the new number of shares. If per-share computations reflect such changes in the number of shares, that fact shall be disclosed.”